OpenStudy (anonymous):

Use multiples to write two fractions equivalent to 9/10

OpenStudy (anonymous):

\(\large \frac{9}{10} \cdot \frac{2}{2}= \) ??? \(\large \frac{9}{10} \cdot \frac{3}{3}= \) ???

OpenStudy (anonymous):

ummm id think thats right...

OpenStudy (anonymous):

carry out the operation.... wat u got?